OpenStudy (anonymous):

Solve 9x2 + 3 = -8x by using the quadratic formula

OpenStudy (anonymous):

\(\large 9x^2+3=-8x \) get this into standard from first to identify your a, b, c \(\large 9x^2+8x+3=0 \) now can you identify a, b, c from the quadratic equation?

OpenStudy (anonymous):

\(\large ax^2+bx+c=0 \) the quadratic formula: \(\large x=\frac{-b\pm \sqrt{b^2-4ac}}{2a} \)